OpenStudy (chriss):

Evaluate \[\int\limits\int\limits_{R}\frac{x+2y}{\cos(x-y)}dA \] where R is the parallelogram bounded by the lines y=x, y=x-1, x+2y=0, and x+2y=2 I could use some help setting this up.

OpenStudy (zarkon):

looks like you will have two double integrals

OpenStudy (chriss):

I've got my four corner points for the parallelogram, but I get so confused setting up my limits of integration, and figuring out what order to integrate in. My corners are (0,0), (2/3,2/3), (2/3,-1/3), (4/3,1/3)

OpenStudy (zarkon):

so one will have x=0 to 2/3 the other will have x=2/3 to 4/3

OpenStudy (chriss):

Two double integrals? one for above the x axis and one for below?

OpenStudy (zarkon):



OpenStudy (chriss):

Ok, I see, so x will be my outside variable for both. But I still don't really get how to get my y-limits for the inside part.

OpenStudy (zarkon):

the bounds for y are the equations of the lines

OpenStudy (zarkon):

for the first one -x/2<y<x

OpenStudy (experimentx):

http://www4c.wolframalpha.com/Calculate/MSP/MSP4451a1a5h962f8ag51700005a138998cah189bh?MSPStoreType=image/gif&s=30&w=386&h=299&cdf=RangeControl \[ \int_{0}^{2/3}\int_{x}^{-2x}f(x)dxdy + \int_{2/3}^{4/3}\int_{x-1}^{(2-x)/2}f(x)dxdy\]

OpenStudy (zarkon):

@experimentX it is x+2y=0, and x+2y=2 not 2x+y=0 and 2x+y=2

OpenStudy (experimentx):

sorry ... i was just giving it a try!!

OpenStudy (zarkon):

trying is good :)

OpenStudy (experimentx):

 I haven't done this kinda question before!!